OpenStudy (anonymous):

What is the distance between points P(1, 7) and Q(9, -8)?

OpenStudy (arindameducationusc):

IS THE ANSWER 17... i can explain why

OpenStudy (arindameducationusc):

https://www.ltcconline.net/greenl/courses/154/factor/circle.htm

OpenStudy (anonymous):

Use the distance formula \[\large \bf d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\]

OpenStudy (anonymous):

\[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (aakashsudhakar):

To find the distance between any two points, you have to manipulate the distance formula, which is given as follows: \[d=\sqrt{(y_{2} - y_{1})^2 + (x_{2} - x_{1})^2}\] For this situation, x2 and y2 would correlate to one set of points while x1 and y1 correlate to another. Be sure to use the same coordinate point for x1 and y1 and the other one for x2 and y2!

OpenStudy (anonymous):

It's another way of writing Pythagoras theorem.

OpenStudy (anonymous):

am I doing something wrong? I keep getting the wrong answer. but im plugging the numbers into the formula

OpenStudy (aakashsudhakar):

Talk us through, what are you getting as your numbers?

OpenStudy (anonymous):

What does the formula look like for you when you plug them in?

OpenStudy (anonymous):

d=\[\sqrt{(1-10)^{2}}+\sqrt({1-7})\]except its a continued square root and the 1-7 equation is also squared